OpenStudy (anonymous):

minimizing area: A 36-in. piece of wire is cut into two pieces. One Piece is used to form a circle while the other is used to form a square. How should the wire be cut so that the sum of the areas is minimal? What is the minimum value?

OpenStudy (anonymous):

 The sum of the areas is\[A(r)=\frac{1}{2}(2\pi)r^2 + (36-2\pi r)^2\]You want this to be a \(\textit{minimum}\). This will happen when the first derivative of A = 0. (though this could also be a maximum. To figure out which, you'd need the second derivative I think).

OpenStudy (anonymous):

what do i do next?

OpenStudy (anonymous):

That should get you a value for r.

OpenStudy (anonymous):

Then you work out the length \(2\pi r\), and the length \((36 - 2\pi r)\)